# Real-Valued Somewhat-Pseudorandom Unitaries

### v1

@@ 1 Introduction
Recently, some partial progress has been made towards a construction of PRU. Namely, several works introduced families of unitaries with polynomial seeds and efficient evaluation, but falling short on pdeudorandomness. Lu, Qin, Song, Yao, and Zhao [${}^{+}$23] introduced the notion of Pseudorandom State Scramblers (PRSS), which are unitaries that are only proven to act pseudorandomly on an arbitrary single input state with arbitrary polynomial multiplicity. Namely, any number of copies of the output on one particular input are pseudorandom. Ananth, Gulati, Kaleoglu, and Lin [AGKL23] introduced the notion of Pseudorandom Isometries (PRI), which are not unitary since their output is longer than their input (and the security of the constructions hinges on this property). They proved security of the PRI property for the case of a single input, a polynomial number of Haar-random inputs, or for an inputs which are a subset of computational basis elements (all with arbitrary polynomial multiplicity). The works of [${}^{+}$23, AGKL23] mention a number of applications for the primitives that they defined, including multi-copy security for quantum public-key encryption.
Interestingly, the constructions in [${}^{+}$23, AGKL23] are real-valued. Namely, the unitary family consists of unitaries with only real values. [^2] In [${}^{+}$23] there is an additional construction which uses complex unitaries, but for the PRSS property, a real valued construction suffices. In contrast, Haug, Bharti, and Koh [HBK23] showed that full PRU security cannot be achieved in this way. This is done by observing that if $U$ is real valued, then $U\otimes U$ acts as identity on the maximally entangled state, whereas this is very far from being the case if $U$ is a random unitary. Therefore, if we consider adversaries that are allowed to make entangled queries to the unitary, then it is impossible to construct real-valued pseudorandom unitaries. Indeed, very recently, and concurrently and independently of our work, Metger, Poremba, Sinha, and Yuen [MPSY24] showed that it is possible to construct non-adaptive PRU that are pseudorandom with respect to any set of inputs that cannot change adaptively throughout the querying process, even when those inputs are entangled with each other and/or the environment. Indeed, their construction is inherently complex-valued. The question, therefore, remains:
@@
In this work, we show that it is possible to achieve stronger security notions than [${}^{+}$23, AGKL23] using an extremely simple construction.
@@ 1.4 Future Directions
Contrary to the constructions in [${}^{+}$23], ours is not scalable (a notion introduced in [BS20]). That is, we consider adversaries whose running time is polynomial in the input size of the unitary $n$. In contrast, in a scalable construction, one specifies separately the parameters for the adversary’s complexity and for the input size. Scalable constructions are usually more involved, and in particularly require more computational depth, than non-scalable ones. It remains an open question to find a scalable version of our construction, or to prove lower bounds in this vein.
@@ Acknowledgments
We thank Omri Shmueli for multiple contributions to the results presented in this work.
@@ 2.1 Notation
We denote $N=2^{n}$ and ${N^{\underline{st}}}=\binom{N}{st}(st)!=N(N-1)\cdots(N-st+1)$. We denote the matrix $\ell_{1}$-norm of a linear operator $A$, $\sup_{\|x\|_{1}=1}\|Ax\|_{1}$, by $\|A\|_{1}$.
@@ 2.4 Almost Invariance Under Haar Unitaries
Let $n,q,\ell\in{\mathbb{N}}$. Suppose $\Phi$ is a quantum channel that is a probabilistic mixture of unitaries on $(\ell+n\cdot q)$ qubits. More precisely, $\Phi(\rho)\coloneqq\mathop{{\mathbb{E}}}_{k\leftarrow\mathcal{D}}\left[(I_{\ell}\otimes V_{k}^{\otimes q})\rho(I_{\ell}\otimes(V_{k}^{\dagger})^{\otimes q})\right]$ where $\mathcal{D}$ a distribution over $\{0,1\}^{*}$, and $V_{k}:{\mathbb{C}}^{2^{n}}\rightarrow{\mathbb{C}}^{2^{n}}$ is a unitary for every $k$ in the support of $D$. Suppose for a $(\ell+n\cdot q)$-qubit state $\rho$, $\Phi(\rho)$ is $\epsilon$-almost invariant under $q$-fold Haar unitary, where $\epsilon$ is a negligible function, then the following holds:
@@ 4.2.1 Focusing on Unique States
be the unique restrictions of $\rho$. We start by showing that $\rho$ is close to its unique restriction.
@@
$\|\rho-\rho^{\star}\|\leq O\left((st)^{2}\epsilon\right)~{}.$ (25)
@@
Note that $(I-\Pi^{\star})\rho(I-\Pi^{\star})$ is positive semi-definite. To prove the claim, it is enough to show that
@@
Recall that ${\cal U}^{\star}_{n,{s}t}$ is the set of all $st$ length vectors with unique entries from ${\{0,1\}^{n}}$, and that $g_{z}=\prod_{j,i}g(z_{j,i})$. From the definitions of $U_{g},U_{\pi}$ and claim 4.2.1, for $c_{1}=\mathop{\mathrm{Tr}}[\Pi^{\star}\rho_{g,\pi}]\geq\frac{1}{1-\epsilon(st)^{2}}$, expressing $\rho^{\star}$ in the standard basis, we get
$\rho^{\star}$ $=c_{1}\cdot\mathop{{\mathbb{E}}}_{g,\pi}\left[\sum_{z,z^{\prime}\in{\cal U}^{\star}_{n,{s}t}}g_{z}g_{z^{\prime}}\prod_{\begin{subarray}{c}j\in[s]\\ i\in[t]\end{subarray}}\alpha_{z(j,i)}^{(j)}{\alpha^{*}}^{{(j)}}_{z^{\prime}(j,i)}|{\pi(z)}\rangle\langle{\pi(z^{\prime})}|\right]$ (33)
$=c_{1}\cdot\mathop{{\mathbb{E}}}_{\pi}\left[\sum_{z,z^{\prime}\in{\cal U}^{\star}_{n,{s}t}}\mathop{{\mathbb{E}}}_{g}[g_{z}g_{z^{\prime}}]\prod_{\begin{subarray}{c}j\in[s]\\ i\in[t]\end{subarray}}\alpha_{z(j,i)}^{(j)}{\alpha^{*}}^{{(j)}}_{z^{\prime}(j,i)}|{\pi(z)}\rangle\langle{\pi(z^{\prime})}|\right]~{}.$ (36)
We now notice that for all $z,z^{\prime}\in{\cal U}^{\star}_{n,{s}t}$ it holds that ${\mathbb{E}}_{g}[g_{x}g_{x^{\prime}}^{*}]$ is equal to $1$ if and only if there exists $\sigma\in S_{st}$ (an outer permutation which permutes the positions of the entries) s.t. $z^{\prime}=\sigma(z)$. Otherwise the expectation is equal to 00. For all $(j,i)\in[s]\times[t]$ we define $\alpha^{({j,i})}=\alpha^{({j})}$ (since we implicitly think of the index $(j,i)$ as pointing to the $j$’th qubit group which consists of a $t$-tensor of $|{{\alpha^{({j})}}}\rangle\langle{{\alpha^{({j})}}}|$). Changing the order of summation by $\pi^{-1}$. We get
@@ 4.2.2 Using Orthogonality to Reach Closeness to an Almost Invariant State
For any permutation $\sigma\in S_{st}$, we consider the associated directed graph $G_{\sigma}$, whose vertex set is $[s]\times[t]$, and there is an edge $(j,i)\to(j^{\prime},i^{\prime})$ if and only if $\sigma((j,i))=(j^{\prime},i^{\prime})$. For all $j\in[s]$, we define the $j$-th vertex-block as the set $\{(j,i)\}_{i\in[t]}$. We sometimes completely associate $\sigma$ with $G_{\sigma}$. We associate each vertex with its outgoing edge. For any vertex $v=(j,i)\in G_{\sigma}$ with the outgoing edge $(j^{\prime},i^{\prime})=\sigma((i,j))$, we denote $j_{v}=j$, $j^{\prime}_{v}=j^{\prime}$, namely the block-source and block-destination of $v$ in the graph. We say that $v$ is a crossing vertex if $j_{v}\neq j_{v}^{\prime}$, and otherwise $v$ is non-crossing. We denote the set of crossing vertices by $\mathsf{cv}_{\sigma}$, and will often omit the subscript when $\sigma$ is clear from the context. Likewise, we denote the set of non-crossing vertices by $\overline{\mathsf{cv}_{\sigma}}$. The block edge pattern of $\sigma$ is the vector $\mathsf{p}_{\sigma}\in{\mathbb{N}}^{[s]\times[s]}$, where $\mathsf{p}_{\sigma}[j,j^{\prime}]$ is the number of crossing vertices from block $j$ to block $j^{\prime}$. We say that two permutations are congruent (with respect to $S_{t}^{s}$) if they have the same block edge pattern. It follows that $\sigma,\sigma^{\prime}$ are congruent, denoted $\sigma\cong\sigma^{\prime}$ if and only if there exist $\sigma_{1},\sigma_{2}\in S_{t}^{s}$ s.t. $\sigma^{\prime}=\sigma_{1}\sigma\sigma_{2}$. We overload the notation and use $\mathsf{p}$ also to denote the congruence class corresponding to this pattern. The number of crossing and non-crossing vertices is thus $\left|{\mathsf{cv}}\right|$, and $\left|{\overline{\mathsf{cv}}}\right|$ respectively (so, $\left|{\mathsf{cv}}\right|+\left|{\overline{\mathsf{cv}}}\right|=st$). We note that $\left|{\mathsf{cv}}\right|$, $\left|{\overline{\mathsf{cv}}}\right|$ only depend on the congruence class $\mathsf{p}$.
@@
Partition the space $z\in{\cal U}^{\star}_{n,{s}t}$ into parts that are invariant under “block-permutations”, i.e. under $S_{t}^{s}$. By definition, each such set contains $(t!)^{s}$ different ${z}$ values (recall that all $z(j,i)$ are unique), and the number of partitions is $\frac{{N^{\underline{st}}}}{(t!)^{s}}$. For each partition $P$, define
@@
We show this by over-counting the set of graphs with a given edge pattern and $\left|{\mathsf{cv}}\right|=k$. We first consider all of the crossing edges, of which there are $k$ by definition. For each such edge, the edge pattern already specifies its source and destination blocks, so we need to choose its specific source and origin vertices within the blocks. There are at most $t^{2}$ options for each edge, and thus at most $t^{2k}$ options in general. Then, for all of the other edges, it just remains to go over each of the $s$ blocks of vertices in the graph, and organize the internal edges in the block. We note that any such arrangement can be completed into a permutation in $S_{t}$, by orienting the outgoing and incoming edges of the block towards each other arbitrarily. Therefore, the number of arrangements in each block is at most $\left|{S_{t}}\right|=t!$. It follows that across all blocks, the total number of internal arrangements in bounded by $(t!)^{s}$. The lemma follows from equations 71 and 72. ∎
@@
It is possible to sum over the $x$ elements as follows. Go over all $v$ at arbitrary order, and for each $v$, let $x(v)$ run over all elements in $\{0,1\}^{n}$ that were not selected in the previous $v$’s, for each such value of $x(v)$ continue to pick value for the next $v$ in the order. In order to analyze this expression, we consider a more general expression as follows:
@@
where the indices $y_{i}$ and $z_{i}$ run over all of $\{0,1\}^{n}$ except for the preceding values of the indices. That is, the $y_{1},\ldots,y_{m}$ values are first chosen to be distinct, and then each $z_{i}$ is chosen in order to be distinct from $y_{1},\ldots,y_{m}$ and all preceding $z_{i}$. The function $\delta(\cdot)$ can be an arbitrary polynomial, and the functions $M_{i}$ are monomials, where $\delta$ and $M_{i}$ can act on the set of their operands and their complex conjugates. We only consider setting where the total degree of $M_{i}$ is even. We note that we can always reorder the $z_{i}$’s without effecting the total value of the expression (maintaining the convention that “later” $z_{i}$’s take all values except those of $y_{i}$’s and “previous” $z_{i}$’s and). We let $d_{i}$ denote the total degree of $M_{i}$. We say that an index $z_{i}$ is loaded if $d_{i}\geq 4$, otherwise we say that it is free (by our convention this means that $d_{i}=2$). As a convention, we always order the summation so that the loaded indices are enumerated on before the free ones. We let $\ell^{\prime}$ denote the number of loaded indices. Furthermore, in our setting, any free term $i$ is going to be of the form $M_{i}(\alpha_{z_{i}})=\alpha^{({j})}_{z_{i}}{\alpha^{*}}^{({j^{\prime}})}_{z_{i}}$, or its complex conjugate, for some $j\neq j^{\prime}$. We define the magnitude of $\tau$ as follows, letting $\delta_{0}$ be the maximal value of $\delta$ over all possible inputs that it can take:
@@
We let $d=\sum_{i\in[\ell^{\prime}]}d_{i}$ denote the total degree of all loaded elements. We now recall that $|{\alpha^{({j})}}\rangle,|{\alpha^{({j^{\prime}})}}\rangle$ are orthogonal for $j\neq j^{\prime}$, and therefore $\sum_{z\in\{0,1\}^{n}}\alpha^{({j})}_{z}{\alpha^{*}}^{({j^{\prime}})}_{z}=0$. It therefore follows that if $\tau$ has any free indices, i.e. by our convention if its $\ell$ index is free, then we have that (up to complex conjugation)
@@
where $\left|{\delta_{\ell}}\right|\leq\epsilon\cdot(st)$ from $\epsilon$ flatness. Each term of the form $\alpha^{({j})}_{z_{i}}{\alpha^{*}}^{({j^{\prime}})}_{z_{i}}$ now “joins” $M_{i}$ and so it either creates a new loaded term, if $z_{i}$ was not previously loaded, decreasing the value of $\mathsf{Mag}{}$ by $\epsilon^{2}N$, or adds $2$ to the degree of a pre-existing loaded term if $z_{i}$ was previously loaded, decreasing the value of $\mathsf{Mag}{}$ by $\epsilon$. Furthermore, multiplying by $\delta_{\ell}$ would decrease the value of the “global” delta by a factor of $\epsilon\cdot(st)$. Therefore, we can write $\tau$ as a sum of $\ell$ terms, each of which conforms with the general form of Eq. (76), but with only $\ell-1$ indices (rather than $\ell$), namely:
@@
We prove this inductively over the value of $\ell-\ell^{\prime}$. For the base case, consider the setting where $\ell^{\prime}=\ell$. Notice that $\sum_{y_{1},\ldots,y_{m}}\prod_{i\in[m]}\left|{\alpha^{({j_{i}})}_{y_{i}}}\right|^{2}\leq 1$. Therefore, $\tau$ is a (sub) convex combination of elements of the form $\delta(\alpha_{y_{1}},\ldots,\alpha_{y_{m}})\sum_{z_{1}}M_{1}(\alpha_{z_{1}})\cdots\sum_{z_{\ell}}M_{\ell}(\alpha_{z_{\ell}})$, that are each bounded in absolute value by $\delta_{0}N^{\ell^{\prime}}\epsilon^{d/2}$, which we show below. It follows that if $\tau$ has no free terms, then $\left|{\tau}\right|\leq\mathsf{Mag}(\tau)$, where $\mathsf{Mag}(\tau)$ is given by Eq. (77). Indeed, each such element is a product of $\delta$, times a product of $\ell^{\prime}$ loaded sums. The total number of summands over all the sums is at most $N^{\ell^{\prime}}$ (as $l^{\prime}=l$). Each element in the sum is a product of $d_{i}$ elements from $\alpha$. By flatness, each element of $\alpha$ has absolute value at most $\sqrt{\epsilon}$, and therefore each element in the sum has absolute value at most $\epsilon^{d_{i}/2}$. We get a value that is bounded by $\delta_{0}N^{\ell^{\prime}}\epsilon^{d/2}$ as required. Now consider the case where $\ell>\ell^{\prime}$. In this case, we can write $\tau=\sum_{i\in[\ell]}\tau_{i}$ as above. We notice that for each $\tau_{i}$, we have $\ell_{i}=\ell-1$, and $\ell^{\prime}_{i}\leq\ell^{\prime}+1$, so $\ell-\ell^{\prime}$ shrinks by at most $2$. Therefore we get the bound

### v2

@@ 1 Introduction
Recently, some partial progress has been made towards a construction of PRU. Namely, several works introduced families of unitaries with polynomial seeds and efficient evaluation, but falling short on pdeudorandomness. Lu, Qin, Song, Yao, and Zhao [LQS+23] introduced the notion of Pseudorandom State Scramblers (PRSS), which are unitaries that are only proven to act pseudorandomly on an arbitrary single input state with arbitrary polynomial multiplicity. Namely, any number of copies of the output on one particular input are pseudorandom. Ananth, Gulati, Kaleoglu, and Lin [AGKL23] introduced the notion of Pseudorandom Isometries (PRI), which are not unitary since their output is longer than their input (and the security of the constructions hinges on this property). They proved security of the PRI property for the case of a single input, a polynomial number of Haar-random inputs, or for an inputs which are a subset of computational basis elements (all with arbitrary polynomial multiplicity). The works of [LQS+23, AGKL23] mention a number of applications for the primitives that they defined, including multi-copy security for quantum public-key encryption.
Interestingly, the constructions in [LQS+23, AGKL23] are real-valued. Namely, the unitary family consists of unitaries with only real values. [^2] In [LQS+23] there is an additional construction which uses complex unitaries, but for the PRSS property, a real valued construction suffices. In contrast, Haug, Bharti, and Koh [HBK23] showed that full PRU security cannot be achieved in this way. This is done by observing that if $U$ is real valued, then $U\otimes U$ acts as identity on the maximally entangled state, whereas this is very far from being the case if $U$ is a random unitary. Therefore, if we consider adversaries that are allowed to make entangled queries to the unitary, then it is impossible to construct real-valued pseudorandom unitaries. Indeed, very recently, and concurrently and independently of our work, Metger, Poremba, Sinha, and Yuen [MPSY24] showed that it is possible to construct non-adaptive PRU that are pseudorandom with respect to any set of inputs that cannot change adaptively throughout the querying process, even when those inputs are entangled with each other and/or the environment. Indeed, their construction is inherently complex-valued. The question, therefore, remains:
@@
In this work, we show that it is possible to achieve stronger security notions than [LQS+23, AGKL23] using an extremely simple construction.
@@ 1.4 Future Directions
Contrary to the constructions in [LQS+23], ours is not scalable (a notion introduced in [BS20]). That is, we consider adversaries whose running time is polynomial in the input size of the unitary $n$. In contrast, in a scalable construction, one specifies separately the parameters for the adversary’s complexity and for the input size. Scalable constructions are usually more involved, and in particularly require more computational depth, than non-scalable ones. It remains an open question to find a scalable version of our construction, or to prove lower bounds in this vein.
@@ Acknowledgments
We thank Omri Shmueli for multiple contributions to the results presented in this work. We would also like to thank Yanglin Hu and Marco Patrick Tomamichel for pointing out a gap in a proof in the previous version of the manuscript.
@@ 2.1 Notation
We denote $N=2^{n}$ and ${N^{\underline{st}}}=\binom{N}{st}(st)!=N(N-1)\cdots(N-st+1)$. We denote the trace norm of $A$ by $\|A\|_{1}=\mathop{\mathrm{Tr}}(\sqrt{AA^{\dagger}})$, which is the sum of the singular values.
@@ 2.4 Almost Invariance Under Haar Unitaries
Let $n,q,\ell\in{\mathbb{N}}$. Suppose $\Phi$ is a quantum channel that is a probabilistic mixture of unitaries on $(\ell+n\cdot q)$ qubits. More precisely, $\Phi(\rho)\coloneqq\mathop{{\mathbb{E}}}_{k\leftarrow\mathcal{D}}\left[(I_{\ell}\otimes V_{k}^{\otimes q})\rho(I_{\ell}\otimes(V_{k}^{\dagger})^{\otimes q})\right]$ where $\mathcal{D}$ a distribution over $\{0,1\}^{*}$, and $V_{k}:{\mathbb{C}}^{2^{n}}\rightarrow{\mathbb{C}}^{2^{n}}$ is a unitary for every $k$ in the support of $D$.
Suppose for a $(\ell+n\cdot q)$-qubit state $\rho$, $\Phi(\rho)$ is $\epsilon$-almost invariant under $q$-fold Haar unitary, where $\epsilon$ is a negligible function, then the following holds:
@@ 4.2.1 Focusing on Unique States
Recall that $g_{z}=\prod_{j,i}g(z_{j,i})$. We notice that for all $z,z^{\prime}\in\{0,1\}^{n\cdot st}$ it holds that ${\mathbb{E}}_{g}[g_{z}g_{z^{\prime}}^{*}]$ is equal to $1$ if and only if the binary type of $z$ and $z^{\prime}$ (that is, the histogram of the entries modulus 2) are equal. Otherwise, the expectation is equal to 00. Expressing $\rho$ in the standard basis we get
$\rho$ $=\mathop{{\mathbb{E}}}_{g,\pi}\left[\sum_{z,z^{\prime}\in\{0,1\}^{n\cdot st}}g_{z}g_{z^{\prime}}\prod_{\begin{subarray}{c}j\in[s]\\ i\in[t]\end{subarray}}\alpha_{z(j,i)}^{(j)}{\alpha^{*}}^{{(j)}}_{z^{\prime}(j,i)}|{\pi(z)}\rangle\langle{\pi(z^{\prime})}|\right]$ (22)
$=\mathop{{\mathbb{E}}}_{\pi}\left[\sum_{z,z^{\prime}\in\{0,1\}^{n\cdot st}}\mathop{{\mathbb{E}}}_{g}[g_{z}g_{z^{\prime}}^{*}]\prod_{\begin{subarray}{c}j\in[s]\\ i\in[t]\end{subarray}}\alpha_{z(j,i)}^{(j)}{\alpha^{*}}^{{(j)}}_{z^{\prime}(j,i)}|{\pi(z)}\rangle\langle{\pi(z^{\prime})}|\right]$ (23)
$=\mathop{{\mathbb{E}}}_{\pi}\left[\sum_{z\in\{0,1\}^{n\cdot st}}\sum_{\begin{subarray}{c}z^{\prime}\in\{0,1\}^{n\cdot st}\\ \mathsf{bintype}(z^{\prime})=\mathsf{bintype}(z)\end{subarray}}\prod_{\begin{subarray}{c}j\in[s]\\ i\in[t]\end{subarray}}\alpha_{z(j,i)}^{(j)}{\alpha^{*}}^{{(j)}}_{z^{\prime}(j,i)}|{\pi(z)}\rangle\langle{\pi(z^{\prime})}|\right]~{},$ (24)
where $\mathsf{bintype}(z)$ is the binary type of $z$.
@@
be the unique restrictions of $\rho$. We show that $\rho$ is close to its unique restriction.
@@
$\|\rho-\rho^{\star}\|_{1}\leq O\left((st)^{2}\epsilon\right)~{}.$ (26)
@@
Notice that $\Pi^{\star}\rho(I-\Pi^{\star})=(I-\Pi^{\star})\rho\Pi^{\star}=0$, as $|{\pi(z)}\rangle$ is in the unique restriction if and only if $\langle{\sigma(\pi(z))}|$ is in the unique restriction too (which occurs if and only if the binary type/histogram has $st$ entries of $1$). It follows that $\rho=\Pi^{\star}\rho\Pi^{\star}+(I-\Pi^{\star})\rho(I-\Pi^{\star})$, and as $\Pi^{\star}\rho\Pi^{\star}$ and $(I-\Pi^{\star})\rho(I-\Pi^{\star})$ are positive semi-definite, it is enough to show that
@@
Recall that ${\cal U}^{\star}_{n,{s}t}$ is the set of all $st$ length vectors with unique entries from ${\{0,1\}^{n}}$. From the definitions of $U_{g},U_{\pi}$ and claim 4.5.1, for $c_{1}=\mathop{\mathrm{Tr}}[\Pi^{\star}\rho_{g,\pi}]\geq\frac{1}{1-\epsilon(st)^{2}}$ we get
$\rho^{\star}$ $=c_{1}\cdot\mathop{{\mathbb{E}}}_{\pi}\left[\sum_{z\in{\cal U}^{\star}_{n,{s}t}}\sum_{\sigma\in S_{st}}\prod_{\begin{subarray}{c}j\in[s]\\ i\in[t]\end{subarray}}\alpha_{z(j,i)}^{(j)}{\alpha^{*}}^{{(j)}}_{\sigma(z)(j,i)}|{\pi(z)}\rangle\langle{\sigma(\pi(z))}|\right]~{}.$ (32)
Notice the sum over $z^{\prime}$ changed to sum over $\sigma\in S_{st}$ (an outer permutation which permutes the positions of the entries) as for $z$ with unique entries it holds that $bintype(z)=bintype(z^{\prime})$ if and only if there exists $\sigma\in S_{st}$ s.t. $z^{\prime}=\sigma(z)$.
For all $(j,i)\in[s]\times[t]$ we define $\alpha^{({j,i})}=\alpha^{({j})}$ (since we implicitly think of the index $(j,i)$ as pointing to the $j$’th qubit group which consists of a $t$-tensor of $|{{\alpha^{({j})}}}\rangle\langle{{\alpha^{({j})}}}|$). Changing the order of summation by $\pi^{-1}$. We get
@@ 4.2.2 Using Orthogonality to Reach Closeness to an Almost Invariant State
For any permutation $\sigma\in S_{st}$, we consider the associated directed graph $G_{\sigma}$, whose vertex set is $[s]\times[t]$, and there is an edge $(j,i)\to(j^{\prime},i^{\prime})$ if and only if $\sigma((j,i))=(j^{\prime},i^{\prime})$. For all $j\in[s]$, we define the $j$-th vertex-block as the set $\{(j,i)\}_{i\in[t]}$. We sometimes completely associate $\sigma$ with $G_{\sigma}$.
We associate each vertex with its outgoing edge. For any vertex $v=(j,i)\in G_{\sigma}$ with the outgoing edge $(j^{\prime},i^{\prime})=\sigma((i,j))$, we denote $j_{v}=j$, $j^{\prime}_{v}=j^{\prime}$, namely the block-source and block-destination of $v$ in the graph. We say that $v$ is a crossing vertex if $j_{v}\neq j_{v}^{\prime}$, and otherwise $v$ is non-crossing. We denote the set of crossing vertices by $\mathsf{cv}_{\sigma}$, and will often omit the subscript when $\sigma$ is clear from the context. Likewise, we denote the set of non-crossing vertices by $\overline{\mathsf{cv}_{\sigma}}$.
The block edge pattern of $\sigma$ is the vector $\mathsf{p}_{\sigma}\in{\mathbb{N}}^{[s]\times[s]}$, where $\mathsf{p}_{\sigma}[j,j^{\prime}]$ is the number of crossing vertices from block $j$ to block $j^{\prime}$. We say that two permutations are congruent (with respect to $S_{t}^{s}$) if they have the same block edge pattern. It follows that $\sigma,\sigma^{\prime}$ are congruent, denoted $\sigma\cong\sigma^{\prime}$ if and only if there exist $\sigma_{1},\sigma_{2}\in S_{t}^{s}$ s.t. $\sigma^{\prime}=\sigma_{1}\sigma\sigma_{2}$. We overload the notation and use $\mathsf{p}$ also to denote the congruence class corresponding to this pattern.
The number of crossing and non-crossing vertices is thus $\left|{\mathsf{cv}}\right|$, and $\left|{\overline{\mathsf{cv}}}\right|$ respectively (so, $\left|{\mathsf{cv}}\right|+\left|{\overline{\mathsf{cv}}}\right|=st$). We note that $\left|{\mathsf{cv}}\right|$, $\left|{\overline{\mathsf{cv}}}\right|$ only depend on the congruence class $\mathsf{p}$.
@@
Partition the space $z\in{\cal U}^{\star}_{n,{s}t}$ into parts that are invariant under “block-permutations”, i.e. under $S_{t}^{s}$. By definition, each such set contains $(t!)^{s}$ different ${z}$ values (recall that all $z(j,i)$ are unique), and the number of partitions is $\frac{{N^{\underline{st}}}}{(t!)^{s}}$.
For each partition $P$, define
@@
We show this by over-counting the set of graphs with a given edge pattern and $\left|{\mathsf{cv}}\right|=k$. We first consider all of the crossing edges, of which there are $k$ by definition. For each such edge, the edge pattern already specifies its source and destination blocks, so we need to choose its specific source and origin vertices within the blocks. There are at most $t^{2}$ options for each edge, and thus at most $t^{2k}$ options in general. Then, for all of the other edges, it just remains to go over each of the $s$ blocks of vertices in the graph, and organize the internal edges in the block. We note that any such arrangement can be completed into a permutation in $S_{t}$, by orienting the outgoing and incoming edges of the block towards each other arbitrarily. Therefore, the number of arrangements in each block is at most $\left|{S_{t}}\right|=t!$. It follows that across all blocks, the total number of internal arrangements in bounded by $(t!)^{s}$. The lemma follows from equations 47 and 48.
@@
It is possible to sum over the $x$ elements as follows. Go over all $v$ at arbitrary order, and for each $v$, let $x(v)$ run over all elements in $\{0,1\}^{n}$ that were not selected in the previous $v$’s, for each such value of $x(v)$ continue to pick value for the next $v$ in the order.
In order to analyze this expression, we consider a more general expression as follows:
@@
where the indices $y_{i}$ and $z_{i}$ run over all of $\{0,1\}^{n}$ except for the preceding values of the indices. That is, the $y_{1},\ldots,y_{m}$ values are first chosen to be distinct, and then each $z_{i}$ is chosen in order to be distinct from $y_{1},\ldots,y_{m}$ and all preceding $z_{i}$.
The function $\delta(\cdot)$ can be an arbitrary polynomial, and the functions $M_{i}$ are monomials, where $\delta$ and $M_{i}$ can act on the set of their operands and their complex conjugates. We only consider setting where the total degree of $M_{i}$ is even. We note that we can always reorder the $z_{i}$’s without effecting the total value of the expression (maintaining the convention that “later” $z_{i}$’s take all values except those of $y_{i}$’s and “previous” $z_{i}$’s and).
We let $d_{i}$ denote the total degree of $M_{i}$. We say that an index $z_{i}$ is loaded if $d_{i}\geq 4$, otherwise we say that it is free (by our convention this means that $d_{i}=2$). As a convention, we always order the summation so that the loaded indices are enumerated on before the free ones. We let $\ell^{\prime}$ denote the number of loaded indices. Furthermore, in our setting, any free term $i$ is going to be of the form $M_{i}(\alpha_{z_{i}})=\alpha^{({j})}_{z_{i}}{\alpha^{*}}^{({j^{\prime}})}_{z_{i}}$, or its complex conjugate, for some $j\neq j^{\prime}$.
We define the magnitude of $\tau$ as follows, letting $\delta_{0}$ be the maximal value of $\delta$ over all possible inputs that it can take:
@@
We let $d=\sum_{i\in[\ell^{\prime}]}d_{i}$ denote the total degree of all loaded elements.
We now recall that $|{\alpha^{({j})}}\rangle,|{\alpha^{({j^{\prime}})}}\rangle$ are orthogonal for $j\neq j^{\prime}$, and therefore $\sum_{z\in\{0,1\}^{n}}\alpha^{({j})}_{z}{\alpha^{*}}^{({j^{\prime}})}_{z}=0$. It therefore follows that if $\tau$ has any free indices, i.e. by our convention if its $\ell$ index is free, then we have that (up to complex conjugation)
@@
where $\left|{\delta_{\ell}}\right|\leq\epsilon\cdot(st)$ from $\epsilon$ flatness.
Each term of the form $\alpha^{({j})}_{z_{i}}{\alpha^{*}}^{({j^{\prime}})}_{z_{i}}$ now “joins” $M_{i}$ and so it either creates a new loaded term, if $z_{i}$ was not previously loaded, decreasing the value of $\mathsf{Mag}{}$ by $\epsilon^{2}N$, or adds $2$ to the degree of a pre-existing loaded term if $z_{i}$ was previously loaded, decreasing the value of $\mathsf{Mag}{}$ by $\epsilon$. Furthermore, multiplying by $\delta_{\ell}$ would decrease the value of the “global” delta by a factor of $\epsilon\cdot(st)$. Therefore, we can write $\tau$ as a sum of $\ell$ terms, each of which conforms with the general form of Eq. (52), but with only $\ell-1$ indices (rather than $\ell$), namely:
@@
We prove this inductively over the value of $\ell-\ell^{\prime}$. For the base case, consider the setting where $\ell^{\prime}=\ell$. Notice that $\sum_{y_{1},\ldots,y_{m}}\prod_{i\in[m]}\left|{\alpha^{({j_{i}})}_{y_{i}}}\right|^{2}\leq 1$. Therefore, $\tau$ is a (sub) convex combination of elements of the form $\delta(\alpha_{y_{1}},\ldots,\alpha_{y_{m}})\sum_{z_{1}}M_{1}(\alpha_{z_{1}})\cdots\sum_{z_{\ell}}M_{\ell}(\alpha_{z_{\ell}})$, that are each bounded in absolute value by $\delta_{0}N^{\ell^{\prime}}\epsilon^{d/2}$, which we show below. It follows that if $\tau$ has no free terms, then $\left|{\tau}\right|\leq\mathsf{Mag}(\tau)$, where $\mathsf{Mag}(\tau)$ is given by Eq. (53).
Indeed, each such element is a product of $\delta$, times a product of $\ell^{\prime}$ loaded sums. The total number of summands over all the sums is at most $N^{\ell^{\prime}}$ (as $l^{\prime}=l$). Each element in the sum is a product of $d_{i}$ elements from $\alpha$. By flatness, each element of $\alpha$ has absolute value at most $\sqrt{\epsilon}$, and therefore each element in the sum has absolute value at most $\epsilon^{d_{i}/2}$. We get a value that is bounded by $\delta_{0}N^{\ell^{\prime}}\epsilon^{d/2}$ as required.
Now consider the case where $\ell>\ell^{\prime}$. In this case, we can write $\tau=\sum_{i\in[\ell]}\tau_{i}$ as above. We notice that for each $\tau_{i}$, we have $\ell_{i}=\ell-1$, and $\ell^{\prime}_{i}\leq\ell^{\prime}+1$, so $\ell-\ell^{\prime}$ shrinks by at most $2$. Therefore we get the bound